# Hierarchical in-Context Reinforcement Learning with Hindsight Modular Reflections for Planning

### v1

@@ 4 Performance Evaluation
In this section, we will introduce the three environments for evaluation in Section 4.1, and the experiment/simulation setup and comparison plan in Section 4.2. Then, we will discuss the obtained results and ablation study results in Section 4.3 and 4.4, respectively.
4.1 Environments
@@
We conduct experiments on three environments–ALFWorld [26], Webshop [27], and HotpotQA [28]. These environments are not typical machine learning datasets, rather they are interactive environments that generate observations and receive actions in the form of texts. Although these may seem to be decoupled from the robotic planning, we argue that the action space in these environments can be regarded as plans. With a plan, a low-level executor (e.g., a robot hand) can ground this plan to a sequence of control signals (e.g., move the index finger to the left by 1cm). Our work mainly focuses on the planning part where we generate orders/commands for the robot to perform and a low-level algorithm will translate these orders or commands to the robots’ real action spaces. Therefore, the purpose of these environments is only to test the planning ability of the proposed HCRL.
@@
ALFWorld: ALFWorld is a text-based household environment that requires an agent to make decisions over multiple steps to complete a task such as putting a hot apple on the countertop. We follow the setup in Reflexion and run 134 scenarios across six different tasks, including moving a hidden object, cleaning an object, heating or cooling an object, etc. Typically, a task can be decomposed into several sequential sub-tasks.
@@
Webshop: Webshop is an online benchmark environment that tests the agent’s ability to navigate through websites to locate items, select correct options, and purchase items. There are two types of actions–search for an item on the search page and click a button on the page. The agent will need to extract useful information from the observation and choose the correct option according to the instructions in order to succeed.
HotpotQA: HotpotQA is a Wikipedia-based search simulator with 113K question-answer pairs. The agent needs to search, extract information from search results, and combine information from multiple searches to obtain the final answer. There are three kinds of actions–search an entity, look up a keyword in the searched content, and finish the question with an answer.
4.2 Setup and Comparison
To demonstrate the performance of the proposed HCRL, we compare the proposed HCRL with the following frameworks.
  • Retroformer [32]: Retroformer is a gradient-based framework, where the LLM is frozen and is used as the policy, while another smaller LM is trained to provide verbal feedback on the decisions based on the reward. The authors used Low Rank (LoRA) fine-tuning technique [40] to reduce the number of parameters to be fine-tuned. In our comparison, we use $rank=4$ for LoRA.
  • Reflexion [34]: Reflexion adopts ReAct [33] as the bottom-level actor, which can reason about the current situation and generate tractable reasoning traces. At the end of each episode, Reflexion will generate reflections for the episode following a few reflection examples. These reflections mainly focus on the details of potentially wrong actions.
Unless otherwise specified, we use GPT-3.5-turbo as the LLM backend. For Webshop and HotpotQA, we provide two full trajectories of successful execution to the agent as context. For ALFWorld, following Yao et al. [34], we provide two task-specific examples to the agent for each of the six tasks. We also provide examples for the hindsight reflection. Specifically, we provide two examples of the goal-trajectory reflection and one example of the full-trajectory reflection. Note that since Retroformer uses gradient-based update, we do not count its performance at #Epi=5 as in-context learning results. Similarly, since ADaPT retries to execute the task when failed, we do not count its performance at #Epi=1 as Pass@1 results.
4.3 Results and Analysis
Figure 3: Success rate of ALFWorld tasks. The bars on the bottom represent the success rate obtained from the first episode and the ones stacked on top represent the results obtained after four rounds of HMR reflections.
The evaluation results of the three environments are presented in Table 1 and Figure 3. We can observe that HCRL can achieve decent Pass@1 performance (i.e., the performance at the first episode without any experiences in the past) with only ADaPT achieving a slightly better performance in ALFWorld. The reason is that ADaPT tries to approach a problem multiple times, and if it fails in the first attempt, it will decompose the task into sub-tasks and try to approach it by completing the sub-tasks. Although this seems reasonable, a complex task structure might confuse the LLM and complicate the reasoning process. Therefore, HCRL can outperform ADaPT in Webshop.
Method Alfworld Webshop HotpotQA
#Epi=1 #Epi=5 #Epi=1 #Epi=5 #Epi=1 #Epi=5
Retroformer 62% 100% ($\uparrow$38%) 33% 36% ($\uparrow$3%) 34% 53% ($\uparrow$19%)
ADaPT 71% ($d_{max}$=3) 45% ($d_{max}$=4) - -
Reflexion 54% 78% ($\uparrow$24%) 19% 28% ($\uparrow$9%) 35% 47% ($\uparrow$12%)
ExpeL 59% 64% ($\uparrow$5%) 35% 41% ($\uparrow$6%) 28% 39% ($\uparrow$11%)
HCRL (Ours) 67% 87% ($\uparrow$20%) 69% 83% ($\uparrow$14%) 37% 57 ($\uparrow$20%)
Table 1: Success rate in percentage for three datasets–ALFWorld, Webshop, and HotpotQA. Bold results represent the best results in that column. #Epi represents the number of episodes. Note that Retroformer is a gradient-based method, where they fine-tuned the model with LoRA techniques. $\uparrow$ indicates the performance improvement over #Epi=1.
Another observation from the results is that Retroformer can achieve decent performances when #Epi=5, which owes to its gradient-based nature. However, since the gradient is not directly applied to the LLM agent but to the reflection agent, Retroformer is still constrained by the ability of the reflection agent.
We can also observe that HCRL can outperform Reflexion in both Pass@1 and Pass@4 performances, especially in ALFWorld and Webshop, which typically have long decision-making trajectories. This demonstrates the advantage of the hierarchical structure HCRL has. We cannot directly observe which reflection approach is better, but we will present more insights on this in Section 4.4.
4.4 Ablation Study
To quantize the contribution to the performance of each part of HCRL, we perform an ablation study on the three datasets. We compare HCRL with three different HCRL variants.
  • HCRL-Retry: We provide past failed full trajectories to the agent without any summarization or reflection. Since the trajectories are long, HCRL-Retry will terminate after a few episodes because of exceeding the maximum sequence length of the LLM.
  • HCRL-Reflexion: We adopt Reflexion as the reflection technique where the reflection is on the full trajectory.
  • HCRL w/o Tag: We let the LLM decide what should the next step be instead of guiding it with tags. In this way, besides generating actions in the environment’s action spaces, it can also generate three types of actions–think, propose a goal, and finish a goal. Consequently, the reflection examples are redesigned to remove the tags.
Note that, in the experiment, HCRL-Retry, HCRL-Reflexion, and HCRL-HMR have the same first episode because they all have HCRL as the decision maker and the difference among them is the reflection technique. After we obtain the trajectory of the first episode using HCRL, we directly start from the second episode but with different long-term memories.
@@ 4.4 Ablation Study
Figure 4: Success rate over five episodes on three datasets/environments using GPT-3.5-turbo.
The results are presented in Figure 4. We can observe that the HCRL w/o Tag does not achieve comparable performances with other tag-guided methods, indicating that injecting the prior knowledge of humans’ thought processes is helpful for LLM agents’ decision-making. Moreover, HCRL-Reflexion and HCRL-HMR can outperform HCRL-Retry because experiences in human language can be perceived better by the LLM agent than plain trajectories, not to mention that retry-based methods are not scalable because of the possibility of exceeding the maximum sequence length of the LLMs. Another notable observation is that HCRL-HMR can outperform HCRL-Reflexion, because modular reflections are more efficient in identifying errors agents made, especially in long trajectories and text-heavy environments like ALFWorld (Figure 4(a)) and HotpotQA (Figure 4(c)).
We also present the results obtained using different LLM backends in Figure 5. It can be observed the GPT-4 can indeed bring significant performance improvement for simpler environments like ALFWorld and Webshop. The reason for this improvement is that GPT-4 has better knowledge about the world because it is trained on more and newer data than GPT-3.5. In other words, tasks in ALFWorld, a household environment, and Webshop, an online shopping environment, can be completed with general human skills and knowledge, while questions in HotpotQA require the agent to search for the information. Furthermore, we can observe that GPT-3.5 with GPT-4 HMR can outperform GPT-3.5, which further demonstrates the importance of reflection. Another interesting finding is that GPT-3.5 with GPT-4 HMR can outperform GPT-4 in HotpotQA, indicating that better LLMs do not necessarily lead to better performances in decision-making tasks.
@@
Figure 5: Success rates in percentage obtained with different LLMs. GPT-3.5 and GPT-4 indicate the decision-making and reflection are performed with the same type of LLMs, while GPT-3.5 with GPT-4 HMR indicates GPT-3.5 is used for decision-making while GPT-4 is used for reflection.
5 Limitation and Discussion
@@
The proposed HCRL is essentially a deterministic policy for decision-making, with HMR being the policy gradient update. Though it can achieve decent performance and enhance interpretability, it still underperforms trained large models and has performance bottlenecks in complex tasks such as HotpotQA. The reason for this is that LLMs heavily rely on the knowledge base acquired during their training on a large amount of human language corpus. However, this knowledge base might be biased, causing the LLMs to be stubborn in certain scenarios. This stubbornness is rooted deep in the knowledge of the LLMs. To tackle this, we will work on better reflection techniques or methods beyond reflections in the future to conquer this issue.
6 Conclusion
@@
In this paper, we introduced Hierarchical in-Context Reinforcement Learning (HCRL), an in-context learning framework that decomposes complex tasks into simpler sub-tasks. A high-level policy is responsible for proposing goals that define sub-tasks, while a low-level policy makes decisions to achieve the goals. To enable the framework to improve over multiple episodes, we proposed Hindsight Modular Reflection (HMR), where we introduced low-level reflection and high-level reflection. The proposed framework was evaluated in three environments and the results showed that HCRL can outperform existing frameworks in both Pass@1 performances and multi-episode performances. In the future, we plan to embed HCRL on drones as a planner, with a trained executor to map the plans to robot control signals and evaluate the system’s performance as a whole.
Acknowledgments
If a paper is accepted, the final camera-ready version will (and probably should) include acknowledgments. All acknowledgments go at the end of the paper, including thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.

### v2

@@ IV Performance Evaluation
In this section, we will introduce the three environments for evaluation in Sect. IV-A, and the experiment/simulation setup and comparison plan in Sect. IV-B. Then we will discuss the results obtained and the results of the ablation study in Sect. IV-C and IV-D, respectively.
IV-A Environments
@@
We conducted experiments in three environments–ALFWorld [27], Webshop [28], and HotpotQA [29]. These environments are not typical machine learning datasets; rather, they are interactive environments that generate observations and receive actions in the form of texts. Although these may seem to be decoupled from robotic planning, we argue that the action space in these environments can be regarded as plans. With a plan, a low-level executor (e.g., a robot hand) can ground this plan to a sequence of control signals (e.g., move the index finger to the left by 1cm). Our work mainly focuses on the planning part, where we generate orders/commands for the robot to perform, and a low-level algorithm will translate these orders or commands to the robots’ real action spaces. Therefore, the purpose of these environments is only to test the planning capacity of the proposed RAHL.
@@
ALFWorld: ALFWorld is a household environment that requires an agent to make decisions over multiple steps to complete a task, such as putting a hot apple on the countertop. We follow the setup in Reflexion and run 134 scenarios across six different tasks, including moving a hidden object, cleaning an object, heating or cooling an object, etc. Typically, a task can be decomposed into several sequential sub-tasks.
@@
Webshop: Webshop is an online benchmark environment that tests the agent’s ability to navigate through websites to locate items, select the correct options, and purchase items. There are two types of actions–search for an item on the search page and click a button on the page. The agent will need to extract useful information from the observation and choose the correct option according to the instructions.
@@
Figure 3: Success rate over five episodes in three datasets/environments using GPT-3.5-turbo. The results and confidence intervals are obtained over ten runs.
HotpotQA: HotpotQA is a Wikipedia-based search simulator with 113K question-answer pairs. The agent needs to search, extract information from the search results, and combine information from multiple searches to obtain the final answer. There are three kinds of actions–search an entity, look up a keyword in the searched content, and finish with an answer.
IV-B Setup and Comparison
To show the performance of the proposed RAHL, we compare the proposed RAHL with the following frameworks.
Retroformer [38]: Retroformer is a gradient-based framework, where the LLM is frozen and is used as the policy, while another smaller LM is trained to provide verbal feedback on the decisions based on the reward. The authors used Low Rank (LoRA) fine-tuning [39] to reduce the number of parameters to be fine-tuned. In our comparison, we use $rank=4$ for LoRA.
Reflexion [31]: Reflexion adopts ReAct [30] as the bottom-level actor, which can reason about the current situation and generate tractable reasoning traces. At the end of each episode, Reflexion will generate reflections for the full episode following a few reflection examples.
Unless otherwise specified, we use GPT-3.5-turbo as the LLM back-end. For Webshop and HotpotQA, we provide two full trajectories of successful execution to the agent as context. For ALFWorld, following Yao et al. [31], we provide two task-specific examples to the agent for each of the six tasks. We also provide examples for the hindsight reflection. Specifically, we provide two examples of the goal-trajectory reflection and one example of the full-trajectory reflection. Since ADaPT retries to execute the task when failed, we do not count its performance at #Epi=1 as Pass@1 results.
IV-C Results and Analysis
TABLE I: Success rate in percentage for three datasets–ALFWorld, Webshop, and HotpotQA. The highlighted results represent the best results in that row. #Epi represents the number of episodes.
Datasets #Epi Retroformer ADaPT Reflexion ExpeL RAHL
Alfworld 1 62% 71% 54% 59% 67%
5 100% 78% 64% 87%
Webshop 1 33% 45% 19% 35% 69%
5 36% 28% 41% 83%
HotpotQA 1 34% - 35% 28% 37%
5 53% - 47% 39% 57%
The evaluation results of the three environments are presented in Table I. We can observe that RAHL can achieve decent Pass@1 performance (i.e., the performance at the first episode without any experiences in the past), with only ADaPT achieving a slightly better performance in ALFWorld. The reason is that ADaPT tries to approach a problem multiple times, and if it fails in the first attempt, it will decompose the task into sub-tasks and try to approach it by completing the sub-tasks. Although this seems reasonable, a complex task structure might confuse the LLM and complicate the reasoning process. Therefore, RAHL can outperform ADaPT in the Webshop. Another observation from the results is that Retroformer can achieve decent performances when #Epi=5, which owes to its gradient-based nature. However, since the gradient is not directly applied to the LLM agent but to the reflection agent, the Retroformer is still constrained by the ability of the reflection agent.
@@ IV-C Results and Analysis
Figure 4: Success rates in percentage obtained with different LLMs. GPT-3.5 and GPT-4 indicate the decision-making and reflection are performed with the same type of LLMs, while GPT-3.5 with GPT-4 HMR indicates GPT-3.5 is used for decision-making while GPT-4 is used for reflection. The results and confidence intervals are obtained over ten runs.
We can also observe that RAHL can outperform Reflexion in both Pass@1 and Pass@4 performances, especially in ALFWorld and Webshop, which typically have long decision-making trajectories. This demonstrates the advantage of the hierarchical structure that RAHL has. We cannot directly observe which reflection approach is better, but we will present more insight on this in Sect. IV-D.
Figure 5: (a) The diagram of the system designed for the experiment with Boston Dynamics SPOT. (b) Robot (green spearhead with the head pointing in the camera’s direction) trajectory in the rooms to locate the target person (red dot).
IV-D Ablation Study
To quantify the contribution to the performance of each part of RAHL, we performed an ablation study by comparing RAHL with three different variants of RAHL.
  • RAHL-Retry: We provide past failed full trajectories to the agent without any summarization or reflection. Since the trajectories are long, RAHL-Retry will terminate after a few episodes because of exceeding the maximum sequence length of the LLM.
  • RAHL-Reflexion: We adopt Reflexion as the reflection technique where the reflection is on the full trajectory.
  • RAHL w/o Tag: We let the LLM decide what the next step should be instead of guiding it with tags. In this way, in addition to generating actions in the environment’s action spaces, it can also generate three types of actions: think, propose a goal, and finish a goal. Consequently, the reflection examples have been redesigned to remove the tags.
Note that in the experiment, RAHL-Retry, RAHL-Reflexion, and RAHL-HMR have the same first episode because they all have RAHL as the decision maker, and the difference among them is the reflection technique. After we obtain the trajectory of the first episode using RAHL, we start directly from the second episode but with different long-term memories.
The results are presented in Fig. 3. We can observe that the RAHL w/o Tag does not achieve comparable performances with other tag-guided methods, indicating that injecting the prior knowledge of humans’ thought processes is helpful for LLM agents’ decision-making. Moreover, RAHL-Reflexion and RAHL-HMR can outperform RAHL-Retry because experiences in human language can be perceived better by the LLM agent than plain trajectories, not to mention that retry-based methods are not scalable because of the possibility of exceeding the maximum sequence length of the LLMs. Another notable observation is that RAHL-HMR can outperform RAHL-Reflexion, because modular reflections are more efficient in identifying errors agents made, especially in long trajectories and text-heavy environments like ALFWorld (Fig. 3(a)) and HotpotQA (Fig. 3(c)).
We also present the results obtained using different LLM backends in Fig. 4. It can be observed that GPT-4 can indeed bring about a significant performance improvement for simpler environments such as ALFWorld and Webshop. The reason for this improvement is that GPT-4 has better knowledge about the world because it is trained on more and newer data than GPT-3.5. In other words, tasks in ALFWorld, a household environment, and Webshop, an online shopping environment, can be completed with general human skills and knowledge, while questions in HotpotQA require the agent to search for the information. Furthermore, we can observe that GPT-3.5 with GPT-4 HMR can outperform GPT-3.5, which further demonstrates the importance of reflection. Another interesting finding is that GPT-3.5 with GPT-4 HMR can outperform GPT-4 in HotpotQA, indicating that better LLMs do not necessarily lead to better performances in decision-making tasks.
V Hardware Experiment
Besides the results obtained on the three benchmark environments, we also performed experiments with the Boston Dynamics SPOT robot dog. The main focus of this experiment is to show the ability of the proposed RAHL framework to be implemented in reality. The task in the experiment is navigation, where the robot needs to navigate to a person in the room next to it. To go to the destination room, the robot needs to look for a door and then look for the person after going through the door. The system used in the experiment and the picture of the SPOT robot are shown in Fig. 5(a).
We design two experiments. In the first test, the robot faces the door in the beginning. The goal of this trial is to test the compatibility of GPT-4-turbo as the image analyst and GPT-4o as the decision-maker. As a result, the robot was able to recognize the door and go through the door to look for the person. The second test increases the difficulty by positioning the robot in a corner of the room. To reach the person, one needs to scan the first room to find the door and then go through the door to find the person. The trial focuses more on the ability of the GPT-4o as the decision-maker. The robot was successful in finding the person in the end after exploring the first room. The trajectory of the second trial is presented in Fig. 5(b).
VI Limitation and Discussion
@@
The proposed RAHL is essentially a deterministic policy for decision-making, with HMR being the update of the policy gradient. Though it can achieve decent performance and enhance interpretability, it still underperforms trained large models and has performance bottlenecks in complex tasks such as HotpotQA. The reason for this is that LLMs rely heavily on the knowledge base acquired during their training on a large human language corpus. However, this knowledge base might be biased, causing LLMs to be stubborn in certain scenarios. This stubbornness is deep rooted in the knowledge of LLM. To tackle this, we will work on better reflection techniques or methods beyond reflections in the future to conquer this issue.
VII Conclusion
@@
In this paper, we introduced Retrieval-Augmented in-context reinforcement Learning (RAHL), an in-context learning framework that decomposes complex tasks into simpler sub-tasks. A high-level policy is responsible for proposing goals that define sub-tasks, while a low-level policy makes decisions to achieve the goals. To enable the framework to improve over multiple episodes, we proposed Hindsight Modular Reflection (HMR), where we introduced low-level reflection and high-level reflection. The proposed framework was evaluated in three environments and the results showed that RAHL can outperform existing frameworks in both Pass@1 performances and multi-episode performances. In the future, we plan to embed RAHL on drones as a planner, with a trained executor to map the plans to robot control signals and evaluate the system’s performance as a whole.
Acknowledgement
This work is supported by the NSF RTML Award No. CCF-1937403.